# Date of publication xxxx 00, 0000, date of current version xxxx 00, 0000. xxxx

### v1

@@ Abstract
Fourier Analysis Network (FAN) was recently proposed as a simple way to improve neural network performance by replacing part of ReLU activations with sine and cosine functions. Although several studies have reported small but consistent gains across tasks, the underlying mechanism behind these improvements has remained unclear. In this work, we show that only the sine activation contributes positively to performance, whereas the cosine activation tends to be detrimental. Our analysis reveals that the improvement is not a consequence of the sine function’s periodic nature; instead, it stems from the function’s local behavior near $x=0$, where its non-zero derivative mitigates the vanishing-gradient problem. We further show that FAN primarily alleviates the dying-ReLU problem, in which a neuron consistently receives negative inputs, produces zero gradients, and stops learning. Although modern ReLU-like activations, such as Leaky ReLU, GELU, and Swish, reduce ReLU’s zero-gradient region, they still contain input domains where gradients remain significantly diminished, contributing to slower optimization and hindering rapid convergence. FAN addresses this limitation by introducing a more stable gradient pathway. This analysis shifts the understanding of FAN’s benefits from a spectral interpretation to a concrete analysis of training dynamics, leading to the development of the Dual-Activation Layer (DAL), a more efficient convergence accelerator. We evaluate DAL on three tasks: classification of noisy sinusoidal signals versus pure noise, MNIST digit classification, and ECG-based biometric recognition. In all cases, DAL models converge faster and achieve equal or higher validation accuracy compared to models with conventional activations.
@@ I Introduction
Finally, we apply DAL to real-world tasks — handwritten digit classification (MNIST) and ECG-based biometric recognition (ECG-ID) — and observe consistent acceleration in convergence across both problems. In the ECG-ID task, we observed not only faster convergence but also an improvement in the network’s final performance, perhaps because it is a more complex problem.
@@ II Related works
Beyond monotonic activations, several works have explored the use of periodic or oscillatory functions in neural networks. Periodic activations have been investigated in the context of implicit neural representations where sinusoidal activations are essential for modeling complex signals, as in SIREN [Sitzmann2020SIREN]. Other works have used Fourier features or encodings to enrich input representations and improve generalization [Tancik2020FourierFeatures]. However, these approaches are fundamentally distinct from Fourier Analysis Networks (FAN). They primarily employ sinusoidal transformations on inputs or features for enhanced representational capacity or complex function modeling, whereas FAN’s mechanism relies on mixing sinusoidal and ReLU-like functions as activations to directly improve the network’s training dynamics and overall performance.
The Fourier Analysis Network (FAN) architecture proposed by Dong et al. [Dong2025] introduced a hybrid activation layer composed of ReLU, sine, and cosine functions. FAN demonstrated small but consistent improvements across diverse tasks, including language modeling, time-series forecasting, and symbolic expression learning. Despite these empirical gains, the underlying mechanism behind the improved performance remained unclear. Dong et al. hypothesized that the periodic nature of sine and cosine could help capture periodic patterns in data, though no evidence supporting this explanation was provided.
Subsequent studies have examined the use of FAN across a range of application domains. Dere et al. [Dere2025] incorporated FAN layers into the fully connected stages of an electromyography (EMG) gesture-recognition network (EMG-Net), reporting a reduction in gesture-classification error from 12.22% to 6.92%. Zeng et al. [Zeng2025] replaced the fully connected layers in a passive seismic source-localization framework with FAN layers; however, because their work does not include a clear comparison against baseline architectures, the magnitude of the improvement remains difficult to assess.
@@ V-B ECG-ID Classification
The dataset was partitioned into two stratified folds. For each run, one fold were used for training and the remaining fold was split evenly into validation and test sets. In this work, we report performance based on the validation set, omitting a separate test set as the focus remains on comparative training dynamics.
@@ V-B1 Without batch normalization
Figure 15: Evolution of training and test accuracies and losses over M=500M{=}500 epochs for ECG-ID classification, averaged over N=10N{=}10 runs, without batch normalization. The model using (ReLU, Sine, 1:1) activations converges faster and attains statistically better performance than the pure ReLU baseline.
@@ VI Conclusions
In this work, we investigated the mechanism underlying the performance improvements reported for Fourier Analysis Networks (FAN). Although FAN was originally interpreted as benefiting from the periodic nature of sine and cosine activations, our experiments demonstrate that this explanation is insufficient. By isolating the contribution of each activation function, we showed that only the sine activation contributes positively to performance, while the cosine activation consistently degrades convergence and accuracy.
@@
Based on these insights, we proposed the Dual-Activation Layer (DAL), which combines two activation functions in a single layer — for example, ReLU and sine. Across a variety of tasks, including synthetic 1D signal classification, MNIST digit recognition, and ECG-ID signal classification, DAL consistently accelerates convergence and often reaches final performance slightly better than that of standard activations. The computational overhead introduced by the sine function is minimal, ensuring DAL operates with nearly the same efficiency as a standard ReLU layer.
In summary, we have established that the performance gain attributed to Fourier Analysis Networks (FAN) is not due to the spectral decomposition hypothesis, but rather to the introduction of activation functions that stabilize the gradient flow near the origin, specifically mitigating the “dying ReLU” problem. Our proposed Dual-Activation Layer (DAL) serves as a robust convergence accelerator.
@@
[Uncaptioned image] Sam Jeong received the B.S. degree in electrical engineering with emphasis in electronic systems in 2017, followed by M.S. degree in electrical engineering with area of concentration in biomedical engineering in 2021, both from the University of São Paulo (USP), Brazil. He is currently pursuing the Ph.D. degree in electrical engineering with the Department of Electronic Systems, USP. His research interest includes signal and image processing, medical signal and image analysis, and machine learning.
[Uncaptioned image] Hae Yong Kim received the B.S. and M.S. degrees (with distinctions) in computer science and the Ph.D. degree in electrical engineering from the Universidade de São Paulo (USP), Brazil, in 1988, 1992 and 1997, respectively. He is currently an Associate Professor with the Department of Electronic Systems Engineering, USP. He is the author of more than 100 articles and holds three patents. His research interests include image processing, machine learning, medical image processing and computer security. Dr. Kim and colleagues received the 6th edition of the Petrobras Technology Award in the “Refining and Petrochemical Technology” category (2013); the “Best Paper in Image Analysis” award at the Pacific-Rim Symposium on Image and Video Technology (2007); and the Thomson ISI Essential Science Indicators “Hot Paper” award, for writing one of the top 0.1% of the most cited computer science papers (2005).
[…]

### v2

@@ Abstract
Fourier Analysis Network (FAN) was recently proposed as a simple way to improve neural network performance by replacing part of Rectified Linear Unit (ReLU) activations with sine and cosine functions. Although several studies have reported small but consistent gains across tasks, the underlying mechanism behind these improvements has remained unclear. In this work, we show that only the sine activation contributes positively to performance, whereas the cosine activation tends to be detrimental. Our analysis reveals that the improvement is not a consequence of the sine function’s periodic nature; instead, it stems from the function’s local behavior near $x=0$, where its non-zero derivative mitigates the vanishing-gradient problem. We further show that FAN primarily alleviates the dying-ReLU problem, in which a neuron consistently receives negative inputs, produces zero gradients, and stops learning. Although modern ReLU-like activations, such as Leaky ReLU, GELU, and Swish, reduce ReLU’s zero-gradient region, they still contain input domains where gradients remain significantly diminished, contributing to slower optimization and hindering rapid convergence. FAN addresses this limitation by introducing a more stable gradient pathway. This analysis shifts the understanding of FAN’s benefits from a spectral interpretation to a concrete analysis of training dynamics, leading to the development of the Dual-Activation Layer (DAL), a more efficient convergence accelerator. We evaluate DAL on three tasks: classification of noisy sinusoidal signals versus pure noise, MNIST digit classification, and Electrocardiogram (ECG)-based biometric recognition. In all cases, DAL models converge faster and achieve equal or higher validation accuracy compared to models with conventional activations.
@@ I Introduction
Finally, we apply DAL to real-world tasks — handwritten digit classification (MNIST) and electrocardiogram-based biometric recognition (ECG-ID) — and observe consistent acceleration in convergence across both problems. In the ECG-ID task, we observed not only faster convergence but also an improvement in the network’s final performance, perhaps because it is a more complex problem.
@@ II Related works
Beyond monotonic activations, several works have explored the use of periodic or oscillatory functions in neural networks. Periodic activations have been investigated in the context of implicit neural representations where sinusoidal activations are essential for modeling complex signals, as in SIREN [Sitzmann2020SIREN]. Other works have used Fourier features or encodings to enrich input representations and improve generalization [Tancik2020FourierFeatures]. However, these approaches are fundamentally distinct from FAN. They primarily employ sinusoidal transformations on inputs or features for enhanced representational capacity or complex function modeling, whereas FAN’s mechanism relies on mixing sinusoidal and ReLU-like functions as activations to directly improve the network’s training dynamics and overall performance.
FAN architecture proposed by Dong et al. [Dong2025] introduced a hybrid activation layer composed of ReLU, sine, and cosine functions. FAN demonstrated small but consistent improvements across diverse tasks, including language modeling, time-series forecasting, and symbolic expression learning. Despite these empirical gains, the underlying mechanism behind the improved performance remained unclear. Dong et al. hypothesized that the periodic nature of sine and cosine could help capture periodic patterns in data, though no evidence supporting this explanation was provided.
Subsequent studies have examined the use of FAN across a range of application domains. Dere et al. [Dere2025] incorporated FAN layers into the fully connected stages of an electromyography gesture-recognition network (EMG-Net), reporting a reduction in gesture-classification error from 12.22% to 6.92%. Zeng et al. [Zeng2025] replaced the fully connected layers in a passive seismic source-localization framework with FAN layers; however, because their work does not include a clear comparison against baseline architectures, the magnitude of the improvement remains difficult to assess.
@@ V-B ECG-ID Classification
The dataset was partitioned into two stratified folds. For each run, one fold was used for training and the remaining fold was split evenly into validation and test sets. In this work, we report performance based on the validation set, omitting a separate test set as the focus remains on comparative training dynamics.
@@ V-B1 Without batch normalization
Figure 15: Evolution of training and validation accuracies and losses over M=500M{=}500 epochs for ECG-ID classification, averaged over N=10N{=}10 runs, without batch normalization. The model using (ReLU, Sine, 1:1) activations converges faster and attains statistically better performance than the pure ReLU baseline.
@@ VI Conclusions
In this work, we investigated the mechanism underlying the performance improvements reported for FAN. Although FAN was originally interpreted as benefiting from the periodic nature of sine and cosine activations, our experiments demonstrate that this explanation is insufficient. By isolating the contribution of each activation function, we showed that only the sine activation contributes positively to performance, while the cosine activation consistently degrades convergence and accuracy.
@@
Based on these insights, we proposed the DAL, which combines two activation functions in a single layer — for example, ReLU and sine. Across a variety of tasks, including synthetic 1D signal classification, MNIST digit recognition, and ECG-ID signal classification, DAL consistently accelerates convergence and often reaches final performance slightly better than that of standard activations. The computational overhead introduced by the sine function is minimal, ensuring DAL operates with nearly the same efficiency as a standard ReLU layer.
In summary, we have established that the performance gain attributed to FAN is not due to the spectral decomposition hypothesis, but rather to the introduction of activation functions that stabilize the gradient flow near the origin, specifically mitigating the “dying ReLU” problem. Our proposed DAL serves as a robust convergence accelerator.